OpenStudy (anonymous):

X has half life on 70.86 days, Y has half life of 77.27 days. On Jan 1 at 12:00 o'clock, 815g of X was created and 800g of Y was created at the same time. On which day and time will the mass of X be equal to Y.

OpenStudy (dumbcow):

use the half life to get the decay function for each X,Y \[f(t)= M e^{-kt}\] where M is initial mass set f(t) = M/2 for half life and solve for k \[e^{-kt^*} = \frac{1}{2}\] \[k = \frac{\ln 2}{t^*}\] where t* is half life in days

OpenStudy (dumbcow):

subbing k back in \[f(t) = M (2^{-t/t^*})\]

OpenStudy (dumbcow):

setting masses of X,Y equal you get \[815 (2^{-t/70.86}) = 800 (2^{-t/77.27})\]

OpenStudy (dumbcow):

\[t = \frac{\ln (800/815)}{\ln 2}*\frac{1}{(\frac{1}{77.27} - \frac{1}{70.86})}\]